A person spins the pointer on a wheel and is awarded the amount indicated by the pointer. 1/2 of wheel is worth $2.00, 1/4 of wheel is worth $10.00, and 1/4 of wheel is worth $15.00
It costs $8 to play the game.
Determine: The expectation of a person who plays the game.
The fair price to play the game.
Show your work for full credit.

Answer by jim_thompson5910(35256)

Expectation = (1/2)*(2)+(1/4)*(10)+(1/4)*(15) - 8 = 7.25-8 = -0.75


This means that you expect to lose about 75 cents per game


----------------------

Determining Fair price:

Let x = price to play game


Fair price means expectation = 0


Expectation = (1/2)*(2)+(1/4)*(10)+(1/4)*(15) - x = 0

7.25 - x = 0

x = 7.25


Fair price = 7.25 dollars